# AI for Scientific Discovery is a Social Problem

### v3

@@ 3.3 Solution Three: Standardizing and Curating Scientific Data for Broad Reuse
3.4 Solution Four: Building Accessible and Sustainable Infrastructure
@@
Sustainable infrastructure depends on community ownership rather than reliance on commercial platforms or individual institutions. Open source projects like Jupyter, Conda, and the Scientific Python ecosystem illustrate successful community-driven models that serve broad research needs. Scientific AI systems, however, require additional components such as provenance tracking for reproducibility, specialized metadata schemas, and support for long-running computations.
@@
4 Discussion
Our analysis underscores a central challenge for AI in science: widespread adoption alone does not guarantee deeper understanding or more equitable progress. Many barriers are social rather than technical, including undervaluing contributions to data and infrastructure, fragmented interdisciplinary collaboration, and incentive structures that favor narrow advances over broad impact.
Addressing these challenges requires integrating technical improvements with community-centered strategies. Strengthening collaboration and education, structuring upstream benchmarks, standardizing data practices, and building accessible infrastructure can create a foundation for shared, sustainable progress. Realizing AI’s transformative potential depends as much on aligning community priorities and incentives as on advancing algorithms and hardware.
@@ 4 Discussion
The proposed solutions require coordinated funding, long-term institutional commitment, and policy support, particularly for infrastructure initiatives, which benefit from stability beyond the lifespan of individual grants. Achieving widespread adoption also depends on overcoming community inertia and entrenched tool preferences. Success requires technical excellence combined with sustained engagement, trust-building, and clear migration pathways. While the primary audience for this paper is the scientific community, implementing a truly holistic solution also requires coordinated action from policymakers, AI researchers, funders, and other stakeholders.
@@
International collaboration can reduce disparities in computational access and support equitable participation in scientific AI, but must address data sovereignty, technology transfer, and fair governance.
Future work should establish metrics for democratization that go beyond adoption counts to include contributing institutions, geographic distribution, and capacity gains at under-resourced organizations. Such measures will help ensure progress is genuine, widely shared, and aligned with AI’s transformative potential.

### v4

@@ 1 Introduction
Case Study: The Materials Project
The Materials Project demonstrates successful large-scale data democratization in computational materials science. Launched in 2011 with continuous NSF and DOE funding, it now provides open access to computed properties for over 154,000 known and predicted materials [57, 91]. Its impact stems from deliberate attention to usability alongside technical infrastructure.
Critical design choices enabled broad adoption beyond well-resourced institutions. First, the project uses MongoDB for its backend with JSON-based data formats and well-documented RESTful APIs, prioritizing accessibility over domain-specific complexity [89]. Second, extensive documentation and tutorials target both computational researchers and experimentalists. Third, clear data provenance and versioning allow users to trace computational parameters and reproduce results. Fourth, governance mechanisms incorporate community feedback into curation priorities.
The Materials Project now serves over 600,000 registered users globally, with sustained institutional investment spanning over a decade demonstrating that infrastructure democratization requires long-term commitment rather than short-term grants [91]. Its success illustrates that when data infrastructure receives appropriate community governance and continuous support, researchers at under-resourced institutions can participate meaningfully in computational discovery.
Solution Four: Building Accessible and Sustainable Infrastructure
@@
Sustainable infrastructure depends on community ownership rather than reliance on commercial platforms or individual institutions. Open source projects like Jupyter, Conda, and the Scientific Python ecosystem illustrate successful community-driven models that serve broad research needs. Provenance tracking records the origins and transformations of datasets and analyses, essential for reproducible research [56].
@@
Case Study: Open-Source and Federated Infrastructure
Several initiatives demonstrate viable models for community-owned AI infrastructure. The Hugging Face Hub provides free hosting for over 1 million models, datasets, and applications with built-in version control, standardized metadata, and collaborative features [120]. Its success stems from reducing friction: researchers can share artifacts with simple commands, discover relevant work through search, and reproduce results through integrated tools. The platform’s support for diverse modalities (text, images, scientific data) and integration with standard ML frameworks (PyTorch, TensorFlow, JAX) illustrate how infrastructure can serve broad communities without requiring specialized expertise.
The Open Science Grid (OSG) exemplifies sustainable community governance for high-throughput computing. Operated through a consortium model with NSF and DOE support since 2004, OSG provides US-based researchers with access to distributed computing resources across participating institutions [94, 104]. Its federated architecture allows institutions to contribute resources while maintaining local control, addressing both computational access and data sovereignty concerns. The consortium model distributes operational costs and decision-making, creating resilience against single points of failure.
These examples share critical features: sustainable funding mechanisms (commercial revenue or government investment), active user communities shaping development priorities, low barriers to entry with progressive feature disclosure, and explicit attention to onboarding and documentation. However, both depend on stable institutional support that many scientific domains lack, highlighting continued need for policy interventions recognizing infrastructure as essential research capacity rather than discretionary spending.
Discussion
AI systems already provide meaningful value to scientists by accelerating literature synthesis, assisting with data and code workflows, and helping researchers navigate rapidly expanding knowledge domains [97, 79]. Yet these successes underscore the deeper gap between automating scientific tasks and advancing scientific discovery. Scientific reasoning requires mechanistic understanding, experimental grounding, and participation in community norms. These dimensions are not captured by models optimized for textual correlations. Narratives describing AI systems as “autonomous scientists” blur this distinction and risk diverting attention from the structural roots of the barriers scientists encounter, such as misaligned institutional incentives, incompatible data practices embedded in long-standing workflows, and the absence of shared experimental constraints that give scientific results meaning.
Across disciplines, these obstacles persist for structural reasons rather than because individual researchers fail to coordinate. Data standards must compete with entrenched legacy formats that laboratories have little reason or support to abandon [50]. Benchmark fragmentation reflects the lack of neutral organizations with stable funding to steward collective resources [4]. Infrastructural inequity arises from decades of concentrated investment in a small number of well-resourced institutions [3, 111]. The labor required to address these gaps, including data curation, documentation, and infrastructure maintenance, remains undervalued in academic reward structures that prioritize novel results instead of durable contributions [38]. Interdisciplinary collaboration is shaped by similar constraints. Domain experts often lack access to ML training because curricula lag behind research needs, while ML researchers frequently underappreciate domain-specific limitations because research incentives reward methodological novelty rather than deep domain engagement. Even the uneven adoption of new computational tools is influenced by these structural patterns, since institutions with limited compute or staffing face prohibitive entry costs.
These conditions suggest that the transformative potential of AI for science depends on strengthening the sociotechnical systems in which AI tools operate, rather than relying on increasingly autonomous models. Durable and community-governed benchmarks, practical and enforceable data standards, infrastructure aligned with scientific workflows, and training pathways that bridge epistemic and methodological divides are all essential. Importantly, effective solutions must confront the structural causes of current barriers by realigning incentives, providing stable long-term funding, and creating governance mechanisms that can outlast individual projects. AI will accelerate discovery when scientific communities are able to sustain shared resources, harmonized evaluation practices, and collective stewardship. Progress will come from coordinated communities, not from models acting in isolation from the social and institutional dynamics that shape scientific work.
@@ Discussion
This work synthesizes challenges across a wide range of scientific domains—biology, chemistry, materials science, physics, and climate science—to highlight structural patterns. While these themes are broadly shared, the specific constraints and opportunities vary by field. As a result, some recommendations may need adaptation for particular subdisciplines or regulatory environments. In addition, several of the arguments rely on studies and surveys that do not encompass the full diversity of scientific institutions, especially in underrepresented regions. The rapid pace of progress in machine learning also means that some technical limitations discussed here may evolve, even if the underlying social and institutional dynamics persist. Finally, our discussion of governance and incentives is necessarily high-level; implementing these ideas would require detailed policy design and negotiation among stakeholders.
@@
Future research should investigate how these sociotechnical barriers manifest within specific research communities in order to generate actionable guidance that is sensitive to local norms and constraints. Comparative studies of successful and unsuccessful attempts to establish data standards, shared benchmarks, or community-governed infrastructure would clarify which governance models best support long-term coordination [52, 47]. On the technical side, research into models that incorporate mechanistic priors, integrate with simulation and experimental feedback loops, and provide principled uncertainty quantification may help align AI systems more closely with scientific reasoning [80, 88]. Global-scale studies of compute access and capacity building are also needed to ensure that progress in AI for science benefits a broad scientific community rather than deepening existing disparities [111].